OpenStudy (anonymous):

Which of the following is a subset of the set A: {integers divisible by 2 from 2 to 20}

OpenStudy (anonymous):

There are no choices :) I believe that there is supposed to be?

OpenStudy (anonymous):

{24, 26, 28} {4, 6, 8, 10} {integers divisible by 4 from 2 to 40} {0, 2, 4, 6 …}

OpenStudy (anonymous):

Well, it could technically be either C or D, but I am not quite sure which one it is...

OpenStudy (anonymous):

I wish I could help more.

jimthompson5910 (jim_thompson5910):

It's not C because this set includes numbers that go over 20 and it's not D because this set also has numbers that go over 20 (and it goes on forever whereas the given set does not)

OpenStudy (cwrw238):

its definitely not A as the elements are all over 20

OpenStudy (anonymous):

can you help me with this one

OpenStudy (anonymous):

If h(x) = −3x − 10, find h(−3).

jimthompson5910 (jim_thompson5910):

Replace x with -3 and evaluate h(x) = -3x-10 h(-3) = -3(-3)-10

jimthompson5910 (jim_thompson5910):

Note: -3(-3) means "negative 3 times negative 3"

jimthompson5910 (jim_thompson5910):

Make sure you follow PEMDAS correctly

OpenStudy (anonymous):

-19

jimthompson5910 (jim_thompson5910):

what is -3 times -3?

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

so h(-3) = -3(-3)-10 becomes h(-3) = 9-10

OpenStudy (anonymous):

-19

jimthompson5910 (jim_thompson5910):

Think of 9-10 as "you have $9 in your pocket and you buy something that is $10. Are you in debt? If not, how much do you have left? If you are in debt, how much are you in debt?"